OpenStudy (anonymous):

Solve for x and check. Enter only a number. 3^2x=27 HINT: Express 27 as a power of 3.

ganeshie8 (ganeshie8):

use the hint

OpenStudy (anonymous):

to me i think the answer i think its 3 @ganeshie8

ganeshie8 (ganeshie8):

\(3^{2x} = 27\)

ganeshie8 (ganeshie8):

\(3^{2x} = 3^3\)

ganeshie8 (ganeshie8):

since base is equal, exponents are also equal : \(2x = 3\) \(x = 3/2\)

OpenStudy (anonymous):

need help

ganeshie8 (ganeshie8):

yea ?

OpenStudy (anonymous):

can you help me with this please @ganeshie8

ganeshie8 (ganeshie8):

didnt we finish, i gave u the answer also x = 3/2

ganeshie8 (ganeshie8):

what do u need help wid hmm ?

OpenStudy (anonymous):

thats not the answer and i need help with solving this @ganeshie8

ganeshie8 (ganeshie8):

is the question correct ? is 2x really on exponent ?

OpenStudy (anonymous):

this is how the problem looks 

ganeshie8 (ganeshie8):

yea, then 3/2 is the right answer. look at the steps how we solved.

ganeshie8 (ganeshie8):

you can check x = 3/2 is correct or not by pluggin in x = 3/2 in the given equation

ganeshie8 (ganeshie8):